Fermi National Laboratory
 

Accelerator Update for March 7, 2007 - March 9, 2007

 

Wednesday March 7

The day shift began with Operations monitoring the Tevatron (TeV) store 5265 (90.64) and the Recycler stash (189.4E10), with the Antiproton Source (Pbar) stacking (46.3E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 8:09 AM, a power supply expert found a LCW leak on a Pbar supply (V901). Stacking was stopped during this repair.

At 8:35 AM, a Booster RF station (BRF16) tripped off for the first of two times in three minutes. Operators reset the station both times, but an RF expert reported other problems and bypassed the station. BRF19 was brought online in its place.

At 8:44 AM, the MI beam permit tripped off for the first of two times in eight minutes due to beam position problems. It reset.

The Pbar power supply leak was repaired. Stacking resumed at 9:03 AM.

At 9:09 AM, BRF6 tripped off. It reset.

Pbar experts began a reverse proton study at 9:37 AM.

NuMI experts began an 11- batch slipstacking study at 10:20 AM.

At 10:23 AM, a MI RF station (MIRF14) tripped off.

At 10:29 AM, RF experts reported a LCW leak from a MIRF13 bias supply. Experts investigated MIRF13 & 14.

RF experts brought MIRF14 back online at 10:45 AM.

At 10:59 AM, the NuMI beam permit tripped off due to quadrupole power supply’s current drifting low. It reset.

At 11:04 AM, Operations reported that NuMI had tripped off many times due to bad beam positions. It reset.

At 11:05 AM, a MI-10 kicker (KPS1A) tripped off and wouldn’t reset. Operations contacted an MI expert.

RF experts reported at 11:08 AM that they had completed their work on MIRF13.

At 11:32 AM, Operations reported that a Linac quadrupole power supply’s (QPS113) output was occasionally dropping and preventing Linac from delivering beam.

A MI expert reported at 11:35 AM that KPS1A was up and running.

At 11:51 AM, KPS1A tripped off again. Kicker experts were contacted.

Kicker experts reported at 12:18 PM that they had discovered an old limiter that had been left hooked up to KPS1A. When the rep-rate increased due to the NuMI studies the limiter started tripping. The experts pulled out the old timer.

At 12:37 PM, a Linac RF station (LRF4) was inhibited due to tube sparking. An expert brought the tube up slowly.

Pbar experts ended their reverse protons studies at 1:03 PM. Stacking resumed about 15 minutes later.

At 1:27 PM, MI tripped off for the first of five times in about two hours due to bad beam positions. It reset.

The Neutron Therapy Facility (NTF) reported at 1:34 PM that they had finished treating their last patient for the day.

At 1:34 PM, MI tripped off due to bad beam positions for the first of three times in 25 minutes.

A Pbar expert began a period of reduced intensity stacking studies at 2:26 PM.

At 4:35 PM, LRF4 tripped off. It reset.

At 3 PM, MI tripped off due to a bad beam positions. It reset.

At 3:06 PM, BRF6 tripped off. It reset.

At 3:24 PM, MI tripped off due to bad beam positions. Booster tuning had caused these recent MI trips.

Pbar stacking studies ended and Pbar resumed normal stacking at 3:30 PM.

Operations reported at 3:59 PM that the Recycler had stashed antiprotons once during the day shift.

A MI expert reported at 4:29 PM that the MI BLM aborts had been enabled.

A MI expert began a MI-60 barrier bucket study at 5:07 PM.

Operations reported at midnight that the Recycler had stashed four times during the evening shift.

Thursday March 8

The midnight shift began with Operations monitoring store 5265 (14.73E30) and the stash (354.0E10), with Pbar stacking (8.6E10), and with NuMI taking beam.

TeV experts began an end-of-store study at 12:40 AM.

Operations terminated store 5265 at 2:12 AM.

Operations put the TeV through a dry squeeze at 2:29 AM.

MI tripped off at 2:45 AM due to a BLM abort at MI-30. It reset.

MIRF12 tripped off. The stations filaments needed to time back in. It was reset at 2:47 AM.

Operations put the TeV into shot setup at 3:23 AM.

At 4:22 AM, Operations reported that the MI beam permit was dropping during every load of protons into the TeV. Operations lowered the intensity.

Operations established store 5267 at 5:25 AM with an initial luminosity of 265.64E30.

The new MTest experiment (T958) requested beam at 7:29 AM.

Operations reported at 8 AM that the Recycler had stashed two times during the midnight shift.

At 8:27 AM, the MI beam permit tripped off due to the MI-8 collimators many times over a period of three hours. Operator patiently reset this trip each time.

At 8:32 AM, Operations reported that they were having problems getting the MTest beam past the M03 enclosure. Operations contacted experts who found a power supply on (F:FT3SW) that was supposed to be off. The experts established beam to the MTest experiment at 9:05 AM. (MTest – T958: The University of Texas at Arlington (UTA) and the University of Alberta will test two types of Cerenkov light detectors.)

Pbar experts began a reverse proton study at 12:13 PM.

NuMI began an 11-batch slipstacking study at 12:35 PM.

At 1:06 PM, BRF6 tripped off. It reset.

At 1:15 PM, BRF12 tripped off. It reset.

At 1:49 PM, an RF expert changed a module on MIRF14. MI beam was down for about two minutes.

The NuMI studies ended at 3:53 PM.

Operations reported at 3:57 PM that the Recycler had stashed two times during the day shift.

The MTest experimenters reported at 6:02 PM that they were done with the beam for the day.

Operations reported at midnight that the Recycler had stashed two times during the evening shift.

Friday March 9

The midnight shift began with Operations monitoring store 5267 (56.67E30) and the stash (230.6E10), with Pbar stacking (52.6E10), and with NuMI taking beam.

At 4:05 AM, Operations reported that the Linac Klystron Debuncher gradient went unstable. Operators cycled the soft RF inhibit device and the Debuncher resumed normal operations.

At 5:44 AM, a Linac quadrupole power supply (L:QPS316) went unstable. Operators swapped control cards. At 5:55 AM, nearly all the Linac quadrupole power supplies went unstable. Operators found that tank 3 had a bad driver. Operators swapped out the driver and everything resumed normal operations.

Operations began sending beam to MTest at 7:09 AM.

At 7:54 AM, a Booster vacuum pump tripped off and caused BRF13 & 14 to trip off. Everything reset okay.

Operations terminated store 5267 at 7:57 AM.

Operations reported at 7:57 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Friday

The plans for today and the weekend are to stack, stash, and store. At some time during the day on Monday, March 12, contractors will turn off feeder 23 for about two hours due to its proximity to other feeders that they will work on. There will be no store in the TeV during this work, which means that maintenance work will be allowed for about four hours for most of the accelerators.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/12/2007    email Fermilab